Jan purchased a number of pencils for PHP 180.00 from Christian and sold all except 6 at a profit of PHP 2.00 per pencil. With the total amount received, he could buy 30 more items than before. Find the number of items he bought and the cost of each in pesos.
1
Expert's answer
2020-05-08T19:43:34-0400

"x\;=\;number\;purchased\\x\;-\;6\;=\;number\;sold\\2\;PHP\;=\;profit\\180\;PHP\;=\;\cos t\;of\;each\\number\;sold\;\times\;profit\;=\;30\;\times\;\cos t\;of\;each\\(x\;-\;6)\;\times\;2\;=\;30\;\times\;180\\x\;=\;30\times90\;+\;6\;=\;2706"

Sell cost = cost + profit = 182

Answer: 2706 - number purchased, 182 - sell cost